The sum of two numbers is 60 and their difference is 30. 
The equation formed is                           

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

As per the condition given in the question, difference of the two numbers is 30.
So, 60–2x = 30
Transpose, -2x to RHS then it becomes 2x and 30 to LHS it becomes -30,

60–30 = 2x
30 = 2x
2x–30 = 0

Thanks